# Virtual Sensing for Real-Time Degradation Monitoring of Nuclear Systems: Leveraging DeepONet for Enhanced Sensing Coverage for Digital Twin-Enabling Technology

### v1

@@ 4 Conclusion
The data and code used and/or analysed during this study are available from the corresponding author on reasonable request.
@@
Conceptualization: S.A., R.H. and K.K.; methodology: R.H.; Writing - original draft: R.H.; Review & Editing: All authors; Formal analysis: R.H.; Visualization: R.H. and K.K.; Data curation: R.H.; Supervision: S.A.; Project administration: S.A. and D.A.; All authors have read and agreed to the published version of the manuscript.
@@
During the preparation of this work, the author(s) used a Large Language Model (ChatGPT) for language editing and refinement. The AI was employed solely for improving the clarity and readability of the manuscript, and no content was generated or altered by the AI.

### v2

@@ Acknowledgment
The data and code used and/or analyzed during this study are available from the corresponding author on reasonable request.
@@
Conceptualization: S.A., R.H., F.A. and K.K.; methodology: R.H. and F.A.; Writing - original draft: R.H. and F.A.; Review & Editing: All authors; Formal analysis: R.H.; Visualization: R.H. and F.A.; Data curation: R.H. and F.A.; Supervision: S.A.; Project administration: S.A. and D.A.; All authors have read and agreed to the published version of the manuscript.
@@
During the preparation of this work, the author(s) utilized a Large Language Model (ChatGPT) to assist with language editing and refinement. The AI tool was employed strictly to enhance the clarity, grammar, and overall readability of the manuscript, ensuring the effective communication of ideas. It is important to note that no scientific content, original ideas, or conceptual contributions were generated, altered, or influenced by the AI.